THESE ARE FOURTH GRADE MATH QUESTIONS: PLEASE HELP ME I'M PUZZLED ON THIS

1. AT A PARTY, LOU ATE 6 FEWER APPETIZERS THAN SUSAN. CHRIS ATE 8 MORE THAN LOU. MITCHELL ATE 5 FEWER THAN CHRIS. MITCHELL ATE A TOTAL OF 5 APPETIZERS. HOW MANY DID SUSAN HAVE?
2. THE TIGERS FOOTBALL TEAM GAINED 12 YARDS ON THEIR SECOND PLAY. ON THEIR THRID PLAY, THEY LOST 24 YARDS. ON THEIR FOURTH PLAY, THEY GAINED 31 YARDS AND WERE STOPPED AT THEIR 42-YARD LINE. AT WHAT YARD LINE WERE THEY AFTER THE FIRST PLAY?


Answer by vleith(2983)

Let s be the number susan had
L = s - 6
C = L + 8 = (s-6) + 8 = s + 2
M = c -5 = (s+2)-5 = s - 3
M = 5
So M = 5 = S - 3
S = 8
2) Let X be the starting point after the first play
s = x + 12
t = s - 24 = (x+12) - 24 = x - 12
f = t + 31 = (x-12) + 31 = x + 19
After f they were at their own 42. Which was x + 19 more than after the first play.
42 = x + 19
42 - 19 = x
23 = x
Fourth grade is hard